# Hall $\pi$-subgroups and characters of $\pi^{\prime}$-degree

### v1

@@ Abstract.
The first author’s research is funded by: the European Union Next Generation EU, M4C1, CUP B53D23009410006, PRIN 2022 - 2022PSTWLB Group Theory and Applications; and the INDAM-GNSAGA Project CUP E53C24001950001. The second author gratefully acknowledges support from the AMS-Simons Research Enhancement Grant (AWD-000167 AMS)
@@ 1. Introduction
In [GHSV21], together with A. A. Schaeffer Fry and C. Vallejo, we attempted to extend this theorem from a single prime $p$ to a set $\pi$ of two primes. Among other results, it was shown that if $\pi=\{2,p\}$ and $\gcd(|G|,2p)>1$, then $G$ possesses a nontrivial irreducible character of $\pi^{\prime}$-degree whose values are contained in ${\mathbb{Q}}(e^{2\pi i/p})$. (Here, a character $\chi$ is said to have $\pi^{\prime}$-degree if $\chi(1)$ is not divisible by every prime in $\pi$.) Unfortunately, this phenomenon does not hold in general when $\pi$ consists of two odd primes. For example, as pointed out in [GHSV21, Proposition 4.1], the Tits group $^{2}F_{4}(2)^{\prime}$ has no nontrivial irreducible character of $\{3,5\}^{\prime}$-degree with values in ${\mathbb{Q}}(e^{2\pi i/15})$.
@@
We remark that Theorem A can be seen as a generalization of Theorem 1.1. In fact, when $p$ is odd, Theorem 1.1 follows from the case $\pi=\{p\}$ of Theorem A and the first Sylow theorem.
If we assume that $2\in\pi$, then the statement of Theorem A remains true when $|\pi|\leq 2$, as noted above, but fails in general once $|\pi|\geq 3$. For example, if $G$ is any non-abelian simple group and $\pi=\pi(G)$, then the trivial character $1_{G}$ is the only irreducible character of $G$ with $\pi^{\prime}$-degree. There are also counterexamples with $\pi\varsubsetneq\pi(G)$, such as $(G,\pi)=(\textup{{A}}_{7},\{2,3,5\})$. At present, we do not have a conceptual explanation for why the prime $2$ is special in this context.
@@ 2. Non-Abelian Simple Groups
If $|\pi|=1$ the statement is obviously implied by the Sylow theory. For the other implication, assume that $\textup{{A}}_{n}$ admits a Hall $\pi$-subgroup $H$ and suppose for a contradiction that $|\pi|\geq 2$. In particular let $\pi=\{p_{1},p_{2},\ldots,p_{t}\}$ for some odd primes $p_{1}<p_{2}<\cdots<p_{t}$. We proceed by distinguishing two cases, depending on the solvability of $H$.
(i) First assume that $H$ is a solvable group. By [Hal28] we know that $H$ admits a $\{p_{1},p_{2}\}$-Hall subgroup $K$. It follows that $K$ is a solvable $\{p_{1},p_{2}\}$-Hall subgroup of $\textup{{A}}_{n}$. Moreover, since $2\notin\{p_{1},p_{2}\}$ we also have that $K$ is a solvable $\{p_{1},p_{2}\}$-Hall subgroup of the symmetric group $\textup{{S}}_{n}$. Using [Hal56, Theorem A4], we deduce that $|K|$ is even, and therefore $2\in\{p_{1},p_{2}\}\subseteq\pi$. This clearly contradicts our hypothesis.
(ii) We now consider the case where $H$ is a non-solvable Hall $\pi$-subgroup of $\textup{{A}}_{n}$. Since $2\notin\pi$, we have that $H$ is a non-solvable Hall $\pi$-subgroup of $\textup{{S}}_{n}$. This family of subgroups of symmetric groups is completely described in [Tho66]. In particular, either $H=\textup{{S}}_{n}$ or $n$ is a prime number and $H=\textup{{S}}_{n-1}$. In both cases we would have $2\in\pi$, again contradicting our hypothesis. ∎
@@
We now turn to the proof of Theorem 2.1 for simple groups of Lie type $S\neq{}^{2}F_{4}(2)^{\prime}$. By this, we mean groups of the form $S=G/{\mathbf{Z}}(G)$, where $G:={\mathbf{G}}^{F}$ is the group of fixed points of a simple, simply connected algebraic group $\mathbf{G}$} defined over an algebraically closed field of characteristic $\ell$, under a Steinberg endomorphism $F$ of $\mathbf{G}$}.
@@
Let $({\mathbf{G}}^{*},F^{*})$ be the pair dual to $({\mathbf{G}},F)$, and set $G^{*}:={{\mathbf{G}}^{*}}^{F^{*}}$. Note that $S=[G^{*},G^{*}]$. Let $\mathbf{T}$$^{*}$ be an $F^{*}$-stable maximal torus of $\mathbf{G}$$^{*}$ that is dual to $\mathbf{T}$} in the sense of [Car85, Proposition 4.3.1]. Let $\mathbf{B}$$^{*}$ be an $F^{*}$-stable Borel subgroup of $\mathbf{G}$$^{*}$ containing $\mathbf{T}$$^{*}$. Write $T^{*}:={{\mathbf{T}}^{*}}^{F^{*}}$ and $B^{*}:={{\mathbf{B}}^{*}}^{F^{*}}$. If $\mathbf{U}$$^{*}$ denotes the unipotent radical of $\mathbf{B}$$^{*}$ and $U^{*}:={{\mathbf{U}}^{*}}^{F^{*}}$, then $B^{*}=U^{*}T^{*}$.
@@
The first part follows from, for instance, [GM20, Proposition 24.21], [Mal07, Lemma 2.2], and [Hun24, Lemma 5.8]. The latter part is [GHSV21, Lemma 4.2]. ∎
@@
We write $S={\mathrm{PSL}}_{n}^{\epsilon}(q)$, where the superscript $\epsilon=+1$ corresponds to the linear groups and $\epsilon=-1$ to the unitary groups. We use analogous notation for the related groups, for example $G={\mathrm{SL}}_{n}^{\epsilon}(q)$ and $G^{*}={\mathrm{PGL}}_{n}^{\epsilon}(q)$. Recall that $T^{*}$ is a maximal torus contained in a Borel subgroup $B^{*}$ of $G^{*}$.
By Lemma 2.7, we may assume that $\pi$ and $\pi(T^{*})$ share a common prime, say $p$. In the linear case, we have
@@
It is more convenient to first study the characters of $\widetilde{G}:={\mathrm{GL}}_{n}^{\epsilon}(q)$ and then analyze those of $G={\mathrm{SL}}_{n}^{\epsilon}(q)$ as irreducible constituents of restricted characters. Note that $\widetilde{G}$ is self-dual, and we will identify $\widetilde{G}$ with its dual group.
We first consider the case $\epsilon=1$. Let $\delta$ be a generator of the multiplicative group ${\mathbb{F}}_{q}^{\times}$, and define the semisimple element
@@
[…]
$\{\delta^{(q-1)/p},\delta^{(q-1)(p-1)/p},1^{\,n-2}\}=\{\kappa(g)\delta^{(q-1)/p},\kappa(g)\delta^{(q-1)(p-1)/p},\kappa(g)^{\,n-2}\},$
[…]
Now we have
[…]
${\mathbb{Q}}(\psi)\subseteq{\mathbb{Q}}(\chi_{\widetilde{s}})\subseteq{\mathbb{Q}}(\zeta_{{\mathrm{ord}}(\widetilde{s})})={\mathbb{Q}}(\zeta_{p}).$
@@
$\psi(1)=|G^{*}:{\mathbf{C}}_{G^{*}}(s)|_{\ell^{\prime}}\text{ divides }|G^{*}:T^{*}|_{\ell^{\prime}}.$
[…]
@@
We claim that the restriction of $\chi_{\widetilde{s}}$ to $G$ is irreducible. Suppose otherwise, and let $\psi$ be an irreducible constituent of $\chi_{\widetilde{s}}\!\downarrow_{G}$. Since $\widetilde{G}/G$ is cyclic, $\psi$ is not $\widetilde{G}-$invariant. Let $I$ be the inertia group of $\psi$ in $\widetilde{G}$; then $I$ is a proper subgroup of $\widetilde{G}$.
By Clifford theory, the number of irreducible characters of $\widetilde{G}$ lying above $\psi$ equals the number of irreducible characters of $I$ lying above $\psi$, which is $|I:G|$. Since $|I:G|<|\widetilde{G}:G|=q+1$, there exists distinct (linear) characters $\lambda,\mu\in{\mathrm{Irr}}(\widetilde{G}/G)$ such that
$\lambda\chi_{\widetilde{s}}=\mu\chi_{\widetilde{s}}$.
[…]
$\chi_{\widetilde{s}z_{\lambda}}=\chi_{\widetilde{s}z_{\mu}},$
which implies that $\widetilde{s}z_{\lambda}$ and $\widetilde{s}z_{\mu}$ are $\widetilde{G}-$conjugate. This is impossible, since ${\mathrm{ord}}(\widetilde{s})={\mathrm{ord}}(\widetilde{t})=p$ whereas ${\mathrm{ord}}(z_{\lambda}z_{\mu}^{-1})$ is coprime to $p$ (indeed, ${\mathrm{ord}}(z_{\lambda}z_{\mu}^{-1})\mid(q+1)$ and $p\mid(q-1)$). This contradiction proves the claim.
Finally, arguments similar to those in the linear case show that the restriction of $\chi_{\widetilde{s}}$ to $G$ is trivial on ${\mathbf{Z}}(G)$, has $\pi^{\prime}$-degree, and takes values in ${\mathbb{Q}}(\zeta_{p})$. This completes the proof. ∎
@@
By Propositions 2.3 and 2.9, and the classification of finite simple groups, we may assume that $S\neq{}^{2}F_{4}(2)^{\prime}$ is a simple group of Lie type not of type $A$. By Lemma 2.7, there exists at least one prime lying in both $\pi$ and $\pi(T^{*})$. As before, let $p$ be such a prime.
@@
where $\mathcal{O}$ is the set of $F^{*}$-orbits on the simple roots of the root system of $\mathbf{G}$$^{*}$ associated with $\mathbf{B}$$^{*}$ and $\mathbf{T}$$^{*}$, and $q$ is the absolute value of the eigenvalues of $F^{*}$ on the character group of $\mathbf{T}$$^{*}$; see [Car85, p. 74]. A straightforward case-by-case check, using [Atl, Table 5] for the size of the group of diagonal automorphisms and [Car85, §1.19] for the sizes of the $F^{*}$-orbits, reveals that if an odd prime $p$ is a divisor of $|G^{*}:S|$, then indeed $(|G^{*}:S|_{p})^{2}$ divides $|T^{*}|$, and it follows that $p$ also divides $|T^{*}\cap S|$. (The only exception is the case $S={\mathrm{PSU}}_{3}(q)$ with $(q+1)_{3}=3$, but this was already excluded using Proposition 2.9.) Therefore, we may and will assume from now on that $p\in\pi\cap\pi(T^{*}\cap S)$. Consequently, there exists a semisimple element $s\in G^{*}$ such that
@@
As mentioned above, all the semisimple characters in the Lusztig series $\mathcal{E}(G,s)$ indexed by $s$ have the same degree $|G^{*}:{\mathbf{C}}_{G^{*}}(s)|_{\ell^{\prime}}$. This is a $\pi^{\prime}$-number due to the fact that $s\in T^{*}$, as argued in the proof of Proposition 2.9. Moreover, as noted in the proof of [GHSV21, Proposition 4.5], these semisimple characters take integer values on unipotent elements and therefore have field of values contained in ${\mathbb{Q}}(\zeta_{3})$, by [GHSV21, Lemma 4.3]. Finally, since their degrees are coprime to $3$, their restrictions to ${\mathbf{Z}}(G)$ are multiples of the trivial character. This completes the proof. ∎

### v2

@@ Abstract.
The first author’s research is funded by: the European Union Next Generation EU, M4C1, CUP B53D23009410006, PRIN 2022 - 2022PSTWLB Group Theory and Applications; and the INDAM-GNSAGA Project CUP E53C24001950001. The second author gratefully acknowledges support from the AMS–Simons Research Enhancement Grant (AWD-000167 AMS). We thank Gunter Malle for several helpful comments on an earlier version, particularly regarding Proposition 2.11
@@ 1. Introduction
In [GHSV21], together with A. A. Schaeffer Fry and C. Vallejo, we attempted to extend this theorem from a single prime $p$ to a set $\pi$ of two primes. Among other results, it was shown that if $\pi=\{2,p\}$ and $\gcd(|G|,2p)>1$, then $G$ possesses a nontrivial irreducible character of $\pi^{\prime}$-degree whose values are contained in ${\mathbb{Q}}(e^{2\pi i/p})$. (Here, a character $\chi$ is said to have $\pi^{\prime}$-degree if $\chi(1)$ is not divisible by any prime in $\pi$.) Unfortunately, this phenomenon does not hold in general when $\pi$ consists of two odd primes. For example, as pointed out in [GHSV21, Proposition 4.1], the Tits group $^{2}F_{4}(2)^{\prime}$ has no nontrivial irreducible character of $\{3,5\}^{\prime}$-degree with values in ${\mathbb{Q}}(e^{2\pi i/15})$.
@@
Theorem A may be viewed as a generalization of Theorem 1.1. In fact, when $p$ is odd, Theorem 1.1 follows from the case $\pi=\{p\}$ of Theorem A together with the first Sylow theorem.
If we assume that $2\in\pi$, then the conclusion of Theorem A remains true when $|\pi|\leq 2$, as noted above, but fails in general once $|\pi|\geq 3$. For example, if $G$ is any non-abelian simple group and $\pi=\pi(G)$, then the trivial character $1_{G}$ is the only irreducible character of $G$ with $\pi^{\prime}$-degree. There are also counterexamples with $\pi\varsubsetneq\pi(G)$, such as $(G,\pi)=(\textup{{A}}_{7},\{2,3,5\})$. At present, we do not have a conceptual explanation for why the prime $2$ is special in this context.
@@ 2. Non-Abelian Simple Groups
If $|\pi|=1$ the statement is obviously implied by the Sylow theory. For the other implication, assume that $\textup{{A}}_{n}$ admits a Hall $\pi$-subgroup $H$ and suppose for a contradiction that $|\pi|\geq 2$. In particular let $\pi=\{p_{1},p_{2},\ldots,p_{t}\}$ for some odd primes $p_{1}<p_{2}<\cdots<p_{t}$. Notice that $H$ is solvable, by the Feit–Thompson theorem [FT63]. By [Hal28] we then know that $H$ admits a $\{p_{1},p_{2}\}$-Hall subgroup $K$. It follows that $K$ is a solvable $\{p_{1},p_{2}\}$-Hall subgroup of $\textup{{A}}_{n}$. Moreover, since $2\notin\{p_{1},p_{2}\}$ we also have that $K$ is a solvable $\{p_{1},p_{2}\}$-Hall subgroup of the symmetric group $\textup{{S}}_{n}$. Using [Hal56, Theorem A4], we deduce that $|K|$ is even, and therefore $2\in\{p_{1},p_{2}\}\subseteq\pi$. This clearly contradicts our hypothesis. ∎
We note that the use of the Feit–Thompson theorem can be avoided by instead relying on the known classification of non-solvable Hall subgroups of symmetric groups. Suppose that $H$ is a non-solvable Hall $\pi$-subgroup of $\textup{{A}}_{n}$. Since $2\notin\pi$, it follows that $H$ is also a non-solvable Hall $\pi$-subgroup of $\textup{{S}}_{n}$. This family of subgroups of symmetric groups is completely described in [Tho66]. In particular, either $H=\textup{{S}}_{n}$ or $n$ is a prime number and $H=\textup{{S}}_{n-1}$. In both cases, we would have $2\in\pi$, which again contradicts our hypothesis.
@@
We now turn to the proof of Theorem 2.1 for simple groups of Lie type $S\neq{}^{2}F_{4}(2)^{\prime}$. By this, we mean simple groups of the form $S=G/{\mathbf{Z}}(G)$, where $G:={\mathbf{G}}^{F}$ is the group of fixed points of a simple, simply connected algebraic group $\mathbf{G}$} defined over an algebraically closed field of characteristic $\ell$, under a Steinberg endomorphism $F$ of $\mathbf{G}$}.
@@
Let $({\mathbf{G}}^{*},F^{*})$ be the pair dual to $({\mathbf{G}},F)$, and set $G^{*}:={{\mathbf{G}}^{*}}^{F^{*}}$. Note that $S=[G^{*},G^{*}]$, as $S$ is simple. Let $\mathbf{T}$$^{*}$ be an $F^{*}$-stable maximal torus of $\mathbf{G}$$^{*}$ that is dual to $\mathbf{T}$} in the sense of [Car85, Proposition 4.3.1]. Let $\mathbf{B}$$^{*}$ be an $F^{*}$-stable Borel subgroup of $\mathbf{G}$$^{*}$ containing $\mathbf{T}$$^{*}$. Write $T^{*}:={{\mathbf{T}}^{*}}^{F^{*}}$ and $B^{*}:={{\mathbf{B}}^{*}}^{F^{*}}$. If $\mathbf{U}$$^{*}$ denotes the unipotent radical of $\mathbf{B}$$^{*}$ and $U^{*}:={{\mathbf{U}}^{*}}^{F^{*}}$, then $B^{*}=U^{*}T^{*}$.
@@
The first part follows from, for instance, [MT11, Proposition 24.21], [Mal07, Lemma 2.2], and [Hun24, Lemma 5.8]. The latter part is [GHSV21, Lemma 4.2]. ∎
We handle the linear and unitary groups separately. To unify notation, we write $S={\mathrm{PSL}}_{n}^{\epsilon}(q)$, where the superscript $\epsilon=+1$ corresponds to the linear groups and $\epsilon=-1$ to the unitary groups. We use analogous notation for the related groups, for example $G={\mathrm{SL}}_{n}^{\epsilon}(q)$ and $G^{*}={\mathrm{PGL}}_{n}^{\epsilon}(q)$. It is more convenient to first study the characters of $\widetilde{G}:={\mathrm{GL}}_{n}^{\epsilon}(q)$ and then analyze those of $G={\mathrm{SL}}_{n}^{\epsilon}(q)$ as irreducible constituents of restricted characters. Note that $\widetilde{G}$ is self-dual, and we will identify $\widetilde{G}$ with its dual group. Let $\pi:\widetilde{G}\to G^{*}$ be the natural projection from $\widetilde{G}$ to $G^{*}$.
[…]
Lemma 2.9.
The number of irreducible constituents of the restriction of semisimple character $\chi_{\widetilde{s}}$ to $G$ divides $\gcd({\mathrm{ord}}(\widetilde{s}),n,q-\epsilon)$.
Proof.
[…]
$|{\mathrm{Im}}(\kappa)|\text{ divides }\gcd({\mathrm{ord}}(\widetilde{s}),n,q-\epsilon).$
[…]
The proof of Lemma 2.9 also yeilds the following well-known result; however, we only require the first statement.
Lemma 2.10.
Let $s:=\pi(\widetilde{s})\in G^{*}$, where $\widetilde{s}$ is a semisimple element of $\widetilde{G}$. Suppose that the multiset of eigenvalues of $\widetilde{s}$ is not invariant under multiplication by every nontrivial root of unity. Then the centralizer ${\mathbf{C}}_{{\mathbf{G}}^{*}}(s)$ is connected. Moreover, if the multiset of eigenvalues of $\widetilde{s}$ is not invariant under multiplication by every nontrivial element of the subgroup of ${\mathbb{F}}_{q^{2}-1}^{\times}$ of order $q-\epsilon$, then $\chi_{\widetilde{s}}$ restricts irreducibly to $G$.
Proof.
For every $g\in\pi^{-1}\bigl({\mathbf{C}}_{G^{*}}(s)\bigr)$, we have $g\widetilde{s}g^{-1}=\kappa(g)\widetilde{s}$, and hence $g\widetilde{s}g^{-1}$ and $\widetilde{s}$ have the same eigenvalues. Therefore, if the multiset of eigenvalues of $\widetilde{s}$ is not invariant under multiplication by every nontrivial element of the subgroup of ${\mathbb{F}}_{q^{2}-1}^{\times}$ of order $q-\epsilon$, the homomorphism $\kappa$ must have trivial image, which proves the second statement.
[…]
@@
Recall that $T^{*}$ is a maximal torus contained in a Borel subgroup $B^{*}$ of $G^{*}$. By Lemma 2.7, we may assume that $\pi$ and $\pi(T^{*})$ share a common prime, say $p$. In the linear case, we have
@@
We first consider the case $\epsilon=1$. Let $\delta$ be a generator of ${\mathbb{F}}_{q}^{\times}$, and define the semisimple element
@@
Note that the case $(n,p)=(3,3)$ cannot occur. (Otherwise, by Lemma 2.6, the Borel subgroup $B^{*}$, of order $q^{3}(q-1)^{2}$, would contain a Sylow $3$-subgroup of $G^{*}$ whose order $q^{3}(q^{2}-1)(q^{3}-1)$ has larger $3$-part, leading to a contradiction.) It is then easy to check that the multiset of eigenvalues of $\widetilde{s}$ is not invariant under multiplication by every nontrivial root of unity. By Lemma 2.10, it follows that ${\mathbf{C}}_{{\mathbf{G}}^{*}}(s)$ is connected, and we let $\chi_{s}$ be the unique semisimple character in the series $\mathcal{E}(G,s)$. Using Lemma 2.8, we deduce that
${\mathbb{Q}}(\chi_{s})\subseteq{\mathbb{Q}}(\zeta_{{\mathrm{ord}}(s)})={\mathbb{Q}}(\zeta_{p}).$
@@
$\chi_{s}(1)=|G^{*}:{\mathbf{C}}_{G^{*}}(s)|_{\ell^{\prime}}\text{ divides }|G^{*}:T^{*}|_{\ell^{\prime}}.$
As $|G^{*}:T^{*}|_{\ell^{\prime}}=|G^{*}:B^{*}|$ is a $\pi^{\prime}$-number by Lemma 2.6, we conclude that $\chi_{s}(1)$ is also a $\pi^{\prime}$-number. Also, from the fact that $s\in S=[G^{*},G^{*}]$, we know that $\chi_{s}$ is trivial on ${\mathbf{Z}}({G})$. This completes the proof for linear groups.
@@
Note also that ${\mathrm{ord}}(\widetilde{s})={\mathrm{ord}}(\widetilde{t})=p$, which does not divide $q+1$. Lemma 2.9 then implies that $\chi_{\widetilde{s}}$ restricts irreducibly to $G$. Arguments similar to those in the linear case show that this restriction is trivial on ${\mathbf{Z}}(G)$, has $\pi^{\prime}$-degree, and takes values in ${\mathbb{Q}}(\zeta_{p})$. This completes the proof. ∎
@@
By Propositions 2.3 and 2.11, and the classification of finite simple groups, we may assume that $S\neq{}^{2}F_{4}(2)^{\prime}$ is a simple group of Lie type not of type $A$. By Lemma 2.7, there exists at least one prime lying in both $\pi$ and $\pi(T^{*})$. As before, let $p$ be such a prime.
@@
where $\mathcal{O}$ is the set of $F^{*}$-orbits on the simple roots of the root system of $\mathbf{G}$$^{*}$ associated with $\mathbf{B}$$^{*}$ and $\mathbf{T}$$^{*}$, and $q$ is the absolute value of the eigenvalues of $F^{*}$ on the character group of $\mathbf{T}$$^{*}$; see [Car85, p. 74]. A straightforward case-by-case check, using [Atl, Table 5] for the size of the group of diagonal automorphisms and [Car85, §1.19] for the sizes of the $F^{*}$-orbits, reveals that if an odd prime $p$ is a divisor of $|G^{*}:S|$, then indeed $(|G^{*}:S|_{p})^{2}$ divides $|T^{*}|$, and it follows that $p$ also divides $|T^{*}\cap S|$. (The only exception is the case $S={\mathrm{PSU}}_{3}(q)$ with $(q+1)_{3}=3$, but this was already excluded using Proposition 2.11.) Therefore, we may and will assume from now on that
$p\in\pi\cap\pi(T^{*}\cap S).$
Consequently, there exists a semisimple element $s\in G^{*}$ such that
@@
As mentioned above, all the semisimple characters in the Lusztig series $\mathcal{E}(G,s)$ indexed by $s$ have the same degree $|G^{*}:{\mathbf{C}}_{G^{*}}(s)|_{\ell^{\prime}}$. This is a $\pi^{\prime}$-number due to the fact that $s\in T^{*}$, as argued in the proof of Proposition 2.11. Moreover, as noted in the proof of [GHSV21, Proposition 4.5], these semisimple characters take integer values on unipotent elements and therefore have field of values contained in ${\mathbb{Q}}(\zeta_{3})$, by [GHSV21, Lemma 4.3]. Finally, since their degrees are coprime to $3$, their restrictions to ${\mathbf{Z}}(G)$ are multiples of the trivial character. (Let $\chi$ be such a character and let $\chi_{{\mathbf{Z}}(G)}=\chi(1)\alpha$ for some $\alpha\in{\mathrm{Irr}}({\mathbf{Z}}(G))$. Then $\mathbf{1}_{{\mathbf{Z}}(G)}=\det(\chi)_{{\mathbf{Z}}(G)}=\det(\chi_{{\mathbf{Z}}(G)})=\det(\chi(1)\alpha)=\alpha^{\chi(1)}$, which implies that the order of $\alpha$ divides $(\chi(1),|{\mathbf{Z}}(G)|)$.) This completes the proof. ∎